Monty Company purchased equipment for $237,300 on October 1, 2017. It is estimated that the equipment will have a useful life of 8 years and a salvage value of $13,560. Estimated production is 39,600 units and estimated working hours are 20,300. During 2017, Monty uses the equipment for 530 hours and the equipment produces 1,000 units.Monty Company purchased equipment for $237,300 on October 1, 2017. It is estimated that the equipment will have a useful life
Answer #1

Calculate depreciation :

a)Straight line method for 2017 = (237300-13560/8)*3/12 = $6992

b) Activity method (unit of output) = (237300-13560/39600)*1000 = $5650

c) Activity method (working hours) = (237300-13560/20300)*530 = 5841

d) Sum of year digit = (237300-13560)*7/36*9/12+(237300-13560)*6/36*3/12 = $41951

e) Double decline balance = 237300*25%*9/12+237300*75%*25%*3/12 = $55617